A train is traveling at a constant speed. The table shows the distance the train travels at different times during the trip.
How fast is the train traveling?
_____ mph

Train’s Speed
Hours 3 5 7 9
Miles 495 825 1155 1485

Answer by stanbon(75887)

A train is traveling at a constant speed. The table shows the distance the train travels at different times during the trip.
How fast is the train traveling?
_____ mph

Train’s Speed
Hours 3 5 7 9
Miles 495 825 1155 1485
------------------------
In 2 hrs the distance is (825-495) = 330 miles
rate = dist/time = 330/2 = 165 mph
----------------
Cheers,
Stan H.
----------------